# GFreeDet: Exploiting Gaussian Splatting and Foundation Models for Model-free Unseen Object Detection in the BOP Challenge 2024

### v1

@@ Abstract
In this report, we provide the technical details of the submitted method GFreeDet, which exploits Gaussian splatting and vision Foundation models for the model-free unseen object Detection track in the BOP 2024 Challenge.
††footnotetext: ∗ Equal contribution.††footnotetext: † Corresponding authors.1 Method and Technical Details
@@ 1.1 Problem Statement
Model-free unseen object detection 111bop.felk.cvut.cz/tasks/#ModelFree-2DDet-Unseen is a newly introduced task in the BOP Challenge 2024 222bop.felk.cvut.cz/challenges/bop-challenge-2024. To mitigate the limitation of model-based tasks for open-world scenarios where CAD models are often not available, the model-free setting opts for rapidly learning unseen objects just from video(s). For this purpose, an onboarding stage, i.e., maximum 5 minutes on 1 GPU from reference video(s) showing all possible views at the object, is given for each previously unseen object.
@@ 1.4 Unseen Object Detection
In the matching stage, we aim at assigning each mask proposal an object category and a matching score. Specifically, given the mask proposal , we first compute a global score by averaging the top values of cosine similarity between template global descriptors (), i.e.,
@@ 2.1 Datasets and Evaluation Metrics
The model-free unseen object detection task of BOP Challenge 2024 is evaluated on the BOP-H3 datasets, which consists of three datasets: HOT3D [1], HOPEv2 [14], and HANDAL [3].
@@ 2.2 Results
| MUSE | Unknown | CAD | 42.6 | 47.4 | 27.0 | 39.0 | 1.485 |
| CNOS [10] | FastSAM [18] | CAD | 35.0 | 31.3 | 24.6 | 30.3 | 0.332 |
| CNOS [10] | SAM [6] | CAD | 31.7 | 36.5 | 19.7 | 29.3 | 1.786 |
| GFreeDet (Ours) | FastSAM [18] | Static | 33.8 | 36.4 | 25.5 | 31.9 | 0.278 |
| GFreeDet (Ours) | SAM [6] | Static | 30.9 | 38.4 | 26.4 | 31.9 | 2.136 |
| CNOS [10] | SAM [6] | Static | - | 34.5 | - | - | - |
The results of different unseen object detection methods on the BOP-H3 datasets are presented in Tab. 1. All the results are obtained from the public leaderboard of the BOP Challenge 2024. We compare with all the publicly available methods submitted to both the model-based and model-free unseen object detection tracks until the deadline of the BOP Challenge 2024. For all methods, we report results in terms of the AP (%) on each of the BOP-H3 datasets, the average AP (%) on BOP-H3, and the average processing time per image (s).
As evidenced by the superior results in Tab. 1, the performance of our GFreeDet, utilizing either SAM or FastSAM, stands out in the model-free track for unseen object detection given static onboarding images. Both GFreeDet-SAM and GFreeDet-FastSAM achieve an AP of 31.9 % on the BOP-H3 datasets, with GFreeDet-FastSAM demonstrating a significant speed advantage. In comparison, the prior-art CNOS [10], which only reported results on HOPEv2, falls short in terms of AP when measured against our GFreeDet. Notably, our GFreeDet even surpasses CNOS with CAD-based onboarding in terms of AP, highlighting the great potential of model-free approaches in detecting unseen objects. Despite this, the submitted method MUSE, which employs CAD-based onboarding data, still leads in AP among all the compared methods, although it has yet to disclose its implementation details. Nonetheless, our GFreeDet-FastSAM not only boasts the quickest inference speed among all competitors but also delivers an elegant detection AP, excelling in both aspects regardless of the onboarding type.

### v3

@@ Abstract
We present GFreeDet, an unseen object detection approach that leverages Gaussian splatting and vision Foundation models under model-free setting. Unlike existing methods that rely on predefined CAD templates, GFreeDet reconstructs objects directly from reference videos using Gaussian splatting, enabling robust detection of novel objects without prior 3D models. Evaluated on the BOP-H3 benchmark, GFreeDet achieves comparable performance to CAD-based methods, demonstrating the viability of model-free detection for mixed reality (MR) applications. Notably, GFreeDet won the best overall method and the best fast method awards in the model-free 2D detection track at BOP Challenge 2024. ††footnotetext: ∗ Equal contribution. ††footnotetext: † Corresponding authors.
1 Method and Technical Details
@@ 1.1 Problem Statement
Model-free unseen object detection 111bop.felk.cvut.cz/tasks/#ModelFree-2DDet-Unseen is a newly introduced task in BOP Challenge 2024 [11]222bop.felk.cvut.cz/challenges/bop-challenge-2024. To mitigate the limitation of model-based tasks for open-world scenarios where CAD models are often not available, the model-free setting opts for rapidly learning unseen objects just from video(s). To facilitate this, a short onboarding stage (i.e., a maximum of 5 minutes on 1 GPU) is given to a method for each new object, which involves learning from reference video(s) showing all possible views at the object.
@@ 1.4 Unseen Object Detection
Given the Gaussian model, we render RGB-D synthetic templates covering different orientations of the object using Gaussian splatting as shown in Fig. 2. Here we set to 162 following SAM-6D [8]. For each template , we generate an image-level global descriptor and patch-level descriptors using DINOv2 [12].
@@
In the matching stage, we aim to assign each mask proposal an object category and a matching score. Specifically, given the mask proposal , we first compute a global score by averaging the top values of cosine similarity between template global descriptors (), i.e.,
@@ 2 Experimental Results
| MUSE | Unknown | CAD | 42.6 | 47.4 | 27.0 | 39.0 | 1.485 |
| CNOS [10] | FastSAM [19] | CAD | 35.0 | 31.3 | 24.6 | 30.3 | 0.332 |
| CNOS [10] | SAM [6] | CAD | 31.7 | 36.5 | 19.7 | 29.3 | 1.786 |
| GFreeDet (Ours) | FastSAM [19] | Static | 33.8 | 36.4 | 25.5 | 31.9 | 0.278 |
| GFreeDet (Ours) | SAM [6] | Static | 30.9 | 38.4 | 26.4 | 31.9 | 2.136 |
| CNOS [10] | SAM [6] | Static | - | 34.5 | - | - | - |
@@ 2.1 Datasets and Evaluation Metrics
The model-free unseen object detection task of BOP Challenge 2024 is evaluated on the BOP-H3 datasets, which consist of three datasets: HOT3D [1], HOPEv2 [15], and HANDAL [3].
@@ 2.2 Results
The results of different unseen object detection methods on the BOP-H3 datasets are presented in Tab. 1 and Fig. 3. All the results are obtained from the public leaderboard of the BOP Challenge 2024. We compare with all the publicly available methods submitted to both the model-based and model-free unseen object detection tracks until the deadline of the BOP Challenge 2024. For all methods, we report results in terms of the AP (%) on each of the BOP-H3 datasets, the average AP (%) on BOP-H3, and the average processing time per image (s).
As evidenced by the superior results in Tab. 1, the performance of our GFreeDet, utilizing either SAM or FastSAM, stands out in the model-free track for unseen object detection given static onboarding images. Both GFreeDet-SAM and GFreeDet-FastSAM achieve an AP of 31.9 % on the BOP-H3 datasets, with GFreeDet-FastSAM demonstrating a significant speed advantage. In comparison, the prior-art CNOS [10], which only reported results on HOPEv2, falls short in terms of AP when measured against our GFreeDet. Notably, our GFreeDet even surpasses CNOS with CAD-based onboarding in terms of AP, highlighting the great potential of model-free approaches in detecting unseen objects. Despite this, the submitted method MUSE, which employs CAD-based onboarding data, still leads in AP among all the compared methods, although it has yet to disclose its implementation details. Nonetheless, our GFreeDet-FastSAM not only boasts the quickest inference speed among all competitors but also delivers an elegant detection AP, excelling in both aspects regardless of the onboarding type. Consequently, our GFreeDet-FastSAM has been honored with both the best overall method and the best fast method awards in the model-free 2D detection track at BOP Challenge 2024 [11].
3 Conclusion
In this paper, we have proposed GFreeDet, a model-free unseen object detection method leveraging Gaussian splatting and foundation models. Our unified Gaussian-based object reconstruction and template rendering pipeline enables effective foundation-model-driven segmentation/detection of unseen objects without CAD models. GFreeDet has established a strong baseline for the model-free unseen object detection benchmark on BOP-H3. We anticipate that this approach will inspire further research and applications in mixed reality. However, despite advancements, the current approach, even in its fasted version, is not yet suitable for real-time MR systems. Future works will focus on enhancing the latency and robustness to facilitate real-time applications. Another promising direction would be to further reduce the onboarding cost required by the current setting, making it more accessible and user-friendly.